# Philip ${}^{*\,\sharp}$ Yandex, MIPT &Aleksandr Beznosikov

### v1

@@ 5.2 Fine-tuning experiments
We evaluated the performance of our framework in the context of memory-efficient fine-tuning using the GLUE benchmark (Wang, 2018), a widely-used collection of tasks for evaluating language models. Following the approach from Zhao et al. (2024a), we fine-tuned RoBERTa-base (Liu, 2019) using LoRA (Hu et al., 2021) and GaLore as baselines for comparison. We adhered to the setup described in LoRA, where low-rank updates of rank 8 were applied only to the Q and V matrices. For a detailed description of the experimental setup, see Section A.2.
@@
However, this comparison required a minor modification to FRUGAL compared to the pre-training phase. Instead of selecting active parameters blockwise, we opted for columnwise selection in each matrix. This adjustment was necessary to ensure a fair comparison within a similar memory budget, as the number of trainable parameters in LoRA with rank 8 is approximately $2.5$ times fewer than the number of trainable parameters in any RoBERTa matrix. This transition from blockwise to columnwise selection allowed us to maintain comparable memory usage across methods. For the same reason, we did not include comparisons with BAdam (Luo et al., 2024) in this setup.
@@
As in Section 5.1.1, we conducted additional experiments with FRUGAL using $\rho=0.0$. In this setup, only the classification head is trained using Adam, while the embedding parameters remain frozen, and the remaining parameters are trained using signSGD. The results demonstrate that this training approach barely compromises performance compared to FRUGAL with rank 8, and still outperforms GaLore. Similar to our findings in Section 5.1.1, we observe that the classification head parameters are particularly sensitive to the choice of optimizer, which can be seen in Table 13 where the model’s performance significantly deteriorates when using signSGD for classification head optimization.
@@ 6 Conclusion
In this work, we introduce a new memory-efficient optimization framework, FRUGAL. Within this framework, the optimization space is divided into two subspaces: the first is updated using a state-full algorithm such as Adam, while the second is updated using a state-free algorithm such as signSGD. We prove theoretical convergence guarantees for our framework with SGDM serving as the state-full algorithm and SGD as the state-free algorithm. In experiments involving pre-training and fine-tuning of language models, FRUGAL outperforms other approaches while using the same or smaller memory.
7 Acknowledgments
The work was done in the Laboratory of Federated Learning Problems of the ISP RAS (Supported by Grant App. No. 2 to Agreement No. 075-03-2024-214). The work was partially conducted while Philip Zmushko was a visiting research assistant in Mohamed bin Zayed University of Artificial Intelligence (MBZUAI).
Appendix A Additional experimental details
@@
To begin, we introduce the hyperparameter density $\rho$. This hyperparameter represents the fraction of the total space in Linear layers that is updated with a stateful optimizer. For GaLore, this parameter is equal to $\rho=r/h$, where $r$ is the projection rank, and $h$ is the hidden size of the model. For RandK projection, this parameter can be expressed as $1-s$, where $s$ means sparsity. For BAdam and FRUGAL with the blockwise update, this parameter denotes the ratio of the number of active blocks $a_{\text{block}}$ to the total number of blocks $p$, i.e., $\rho=a_{\text{block}}/p$. When using FRUGAL with the column-wise update, as in Section 5.2, $\rho$ is equal to the ratio of the number of active columns $a_{\text{column}}$ to their total number $h$, i.e., $\rho=a_{\text{column}}/h$.
A.1 Pre-training
@@
Setup. We adopt a LLaMA-based architecture with RMSNorm and SwiGLU (Wang, 2018) activations on the C4 dataset. Following Zhao et al. (2024a), we trained using a batch size of 512 sequences, sequence length of 256, weight decay of 0, and no gradient clipping. We used T5 tokenizer, since it also was trained on C4 with dictionary size equal to 32k. Additionally, the update frequency $T$ is set to 200, and the density $\rho=r/h$ is $0.25$ unless stated otherwise. We also should point out that similarly to Zhao et al. (2024a), we keep Embeddings, RMSNorms (Zhang & Sennrich, 2019), and Logits in the state-full subspace throughout the training and don’t reset the optimizer state for them.
@@
We used standard Adam hyperparameters: $\beta_{1}=0.9,\beta_{2}=0.999,\varepsilon=1e-8$. For all the methods except GaLore, we selected the learning rate equal to the optimal learning rate for Adam, which we determined through a grid search among values $[1e-4,3e-4,1e-3,3e-3]$. FRUGAL’s learning rate for the state-free optimizer was set equal to that for the state-full optimizer for simplicity and ease of tuning. For a fair comparison with GaLore (Zhao et al., 2024a), we conducted experiments with two learning rate values: 1) the one specified by the authors in the original paper, and 2) the optimal learning rate for Adam, as used for other methods. We did this because the learning rate in the original paper could have been optimized for a different number of iterations.
@@
To match the learning rate changes in the first steps of our training with Zhao et al. (2024a), we used a cosine learning rate schedule with restarts, with a warmup of 10% of the steps in a cycle length, and decay of the final learning rate down to 10% of the peak learning rate. To verify that our results are not sensitive to the choice of scheduler, we repeated the experiments for LLaMA-130M with other schedulers. Results for constant with warm-up and cosine (one cycle) with warm-up schedulers can be found in Tables 6 and 6.
@@
Method 100k iterations
Adam 21.88
GaLore, $\rho=0.25$ 24.19
BAdam, $\rho=0.25$ 25.03
FRUGAL, $\rho=0.25$ 23.17
FRUGAL, $\rho=0.0$ 22.64
Table 8: Perplexity of LLaMA-130M models pre-trained on C4 for 20k iterations (2.1B tokens) using SGD and signSGD with different learning rates. $\infty$ means that run diverged. LR stands for learning rate.
SGD signSGD
LR Perplexity LR Perplexity
0.1 184.83 3e-4 40.22
0.3 91.23 1e-3 41.18
1.0 $\infty$ 3e-3 109.32
Method Perplexity
Random 18.50
Ascending 18.54
Descending 18.50
Update frequency $T$ Perplexity
10 18.82
20 18.73
50 18.69
100 18.65
200 18.60
500 18.60
1000 18.61
@@
A.2 Fine-tuning
Setup. The batch size and learning rate values used for FRUGAL in the experiments from Table 4 are presented in Table 12. In all experiments, we set the learning rate for the state-free optimizer to $1/10$ of the learning rate of the state-full optimizer. Other hyperparameters, such as scheduler, number of epochs, maximum sequence length, and warmup ratio, were taken from Hu et al. (2021).
@@ A.2 Fine-tuning
Appendix B Memory estimation

### v2

@@ 6.1 Comparison to existig memory-efficient algorithms
  • Full-rank Training. Training using memory-inefficient AdamW (Loshchilov, 2017). Weights, gradients, and statistics are stored and computed for all parameters. This serves as an upper bound for model performance.
  • BAdam. Luo et al. (2024) proposed a block coordinate descent (BCD)-type optimization method termed BAdam. The parameters are divided into blocks, which are then updated one by one using AdamW. The optimized block is changed every $T$ steps. Although this method was initially proposed only for fine-tuning, it is the closest method to our FRUGAL. Unlike BAdam, in our algorithm, state-free blocks are not frozen but are updated using signSGD.
  • Other Algorithms. Among other relevant methods, ReLoRA (Lialin et al., 2023), MicroAdam (Modoranu et al., 2024), Fira (Chen et al., 2024a), LDAdam (Robert et al., 2024), and Adamem (Vyas et al., 2024) can also be highlighted. However, we did not include them for comparison here for the following reasons: 1. ReLoRA: This method was evaluated in (Zhao et al., 2024a), where it significantly underperformed compared to GaLore. 2. MicroAdam: Its current implementation only supports bfloat16 master weights, whereas our main experiments conducted with mixed precision. 3. Fira, LDAdam, and Adamem: These methods are concurrent works that were published during the final stages of our work. Accordingly, the majority of our experiments were completed before we became aware of these recent developments.
Main results. The results of our experiments are presented in Table 2, which includes both validation perplexity and memory footprint estimations for each method. We compared all memory-efficient methods under the same memory budget with a density $\rho=0.25$. Here, $\rho$ refers to the proportion of Linear layer parameters belonging to the state-full subspace. Similarly to GaLore, non-Linear modules (Embeddings, RMSNorms, Output layer) are optimized with AdamW. See Section A.1 for details.
We conducted a grid search to determine the optimal learning rate for AdamW, which we then applied to FRUGAL and BAdam (Luo et al., 2024). For GaLore (Zhao et al., 2024a), we found that using this same learning rate produced better results than the originally suggested rate. This discrepancy might be attributed to our experiments involving a significantly larger number of training steps than those for which GaLore’s original learning rate was optimized.
Table 2 demonstrates that FRUGAL significantly outperforms memory-efficient baselines across all model sizes with the same memory budget, coming close to the performance of AdamW.
6.2 Zero-density training
Table 2 also reveals a surprising result: FRUGAL with $\rho=0.0$ outperforms both GaLore and BAdam, even when these competing methods use a higher density of $\rho=0.25$. Essentially, for FRUGAL with $\rho=0.0$, the parameters are divided into two parts — a state-full part consisting of the Embeddings, RMSNorms, and Output layer, and a state-free part consisting of all other parameters. This division remains fixed throughout the training. We conducted additional experiments to determine the maximum subset of parameters that can be trained with a state-free optimizer without significant quality degradation. We systematically moved different combinations of the Embeddings, RMSNorms, and Output layer from the state-full to the state-free set and observed the results during the training of LLaMA-130M. Table 4 reveals that the Output layer demonstrates a dramatically higher sensitivity, with changes to its optimizer resulting in severe performance degradation. This finding aligns with results from Zhao et al. (2024b), where the authors demonstrated that most parameters can be trained using SGDM, but the Output layer require training with AdamW.
6.3 LLaMA 3B training
To demonstrate the practical viability of our method for large-scale applications, we evaluated FRUGAL against AdamW on the pre-training of the LLaMA 3B model. Due to computational constraints, we conducted a single training run of 300k steps using a cosine learning rate scheduler with 10% warmup steps. We used a learning rate of 5e-4, weight decay of 0.1, and gradient clipping of 1.0, with other hyperparameters consistent with Section A.1. The results in Table 5 confirm that FRUGAL successfully scales to billion-parameter models without performance degradation, making it a viable option for industrial-scale applications.
6.4 Ablation study
Table 7: Accuracy $\uparrow$ comparison of memory-efficient methods on LLaMA 3.1-8B across 8 commonsense reasoning tasks.
Method Rank BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg
LoRA 32 73.39 88.19 79.48 95.13 83.50 90.24 79.52 84.00 84.18
GaLore 32 74.13 88.36 78.97 94.88 83.82 91.08 80.2 85.80 84.65
FRUGAL 32 74.25 88.47 79.12 95.15 86.11 91.37 79.86 84.80 84.89
FRUGAL 0 71.44 88.47 80.40 94.74 86.35 90.95 81.14 86.00 84.94
We also conducted additional experiments to verify the robustness of our framework to various hyperparameters.
First, we began by evaluating different model architectures. Experiments with GPT-2 124M (Radford et al., 2019) in Table 12 show that FRUGAL maintains its strong advantage over memory-efficient baselines, albeit with a somewhat wider gap to AdamW. Second, an ablation study on the state-full subspace update frequency $T$ in Table 14 shows that the performance keeps improving up to $T=200$. We note that, unlike in Zhao et al. (2024a), the perplexity does not decrease significantly even when reducing the update frequency to $T=10$ ($\sim 0.2$ drop vs. $\sim 4.$ drop for GaLore). A detailed explanation for this result can be found in Appendix D. After that, when using other schedulers, the performance gap between FRUGAL and baselines remains consistent, as shown in Tables 16 and 16. Table 8 shows that the same holds for $\beta_{2}=0.95$ — another popular value for the second moment decay parameter in AdamW-like methods. FRUGAL also provides improvement over baselines for other state-full optimizers, as can be seen in experiments with Lion (Chen et al., 2024b) presented in Table 12. Then, the results of the training in pure bfloat16 are presented in Table 10, demonstrating consistency with our main experiments in Table 2, i.e., FRUGAL significantly outperforms the baselines across these variations. We also conducted experiments to show how perplexity changes with varying $\rho$, and the results are presented in Table 17. Finally, we conducted an experiment to compare different strategies for selecting state-full blocks during training. The results in Table 14 show that there is no significant difference between random and structured block selection.
These experimental results validate that our framework’s superiority is resilient to hyperparameter variations.
7 Fine-tuning experiments
7.1 Fine-tuning RoBERTa on GLUE
@@
We evaluated the performance of our framework in memory-efficient fine-tuning using the GLUE benchmark (Wang, 2018), a widely-used collection of tasks for evaluating language models. Following the approach from Zhao et al. (2024a), we fine-tuned RoBERTa-base (Liu, 2019) using LoRA (Hu et al., 2021) and GaLore as baselines for comparison. We adhered to the setup described in LoRA, where low-rank updates of rank 8 were applied only to the Q and V matrices. See detailed description in Section A.2.
@@
For this experiment we opted for columnwise selection of active parameters. This transition from blockwise to columnwise selection was necessary to maintain comparable memory usage across methods, as the number of trainable parameters in LoRA with rank 8 is approximately $2.5$ times fewer than the number of parameters in any RoBERTa matrix. For the same reason, we did not include comparisons with BAdam (Luo et al., 2024) in this setup.
@@
As in Section 6.1, we conducted additional experiments with FRUGAL using $\rho=0.0$. In this setup, only the classification head is trained using AdamW, while the embedding parameters remain frozen, and the remaining parameters are trained using signSGD. The results demonstrate that this training approach barely compromises performance compared to FRUGAL with rank 8, and still outperforms GaLore.
Similar to our findings in Section 6.1, we observe that the classification head parameters are particularly sensitive to the choice of optimizer, which can be seen in Table 19 where the model’s performance significantly deteriorates when using signSGD for classification head optimization.
7.2 Fine-tuning LLaMA on commonsense reasoning
In addition to our experiments with RoBERTa, we conducted experiments on LLM fine-tuning to evaluate our framework’s performance in this practically important area. To assess this capability, we chose LLaMA 3.1-8B (Dubey et al., 2024) and commonsense reasoning benchmark, as this domain represents a fundamental capability requiring both factual knowledge and logical inference. This benchmark includes 8 subtasks each containing its own training and test splits. Following the setup from Hu et al. (2023) we train the model on a single combined Commonsense170K dataset (Hu et al., 2023). We refer readers to the original paper for detailed descriptions of these tasks and the construction of the Commonsense170K dataset.
Following the experimental protocol from Hu et al. (2023), we apply memory-efficient methods to the same parameter subsets: the Q, K, V, Up, and Down projection matrices. We used the same hyperparameter configuration as in the original work, except for the learning rate, which we varied across [5e-6, 1e-5, 2e-5, 5e-5, 1e-4, 2e-4] for each algorithm to ensure optimal performance. The results in Table 7 show that FRUGAL again slightly outperforms both LoRA and GaLore in the average accuracy across all 8 tasks. Remarkably, our method attains this advantage even with $\rho=0$ (effectively signSGD), requiring zero memory for optimizer state storage while delivering better accuracy.
@@ 8 Conclusion
In this work, we introduce a new memory-efficient optimization framework, FRUGAL. Within this framework, the optimization space is divided into two subspaces: the first is updated using a state-full algorithm such as Adam, while the second is updated using a state-free algorithm such as signSGD. We prove theoretical convergence guarantees for our framework with SGDM serving as the state-full algorithm and SGD as the state-free algorithm. In experiments involving pre-training and fine-tuning of language models, FRUGAL outperforms other approaches.
Acknowledgements
The work on the final version was conducted at Moscow Institute of Physics and Technology and was supported by the Ministry of Economic Development of the Russian Federation (code 25-139-66879-1-0003). The work was partially conducted while Philip Zmushko was a visiting research assistant in Mohamed bin Zayed University of Artificial Intelligence (MBZUAI).
Impact Statement
Our work makes large-scale model training more accessible to the broader research community by reducing memory requirements. This reduction in computational demands not only decreases the financial cost of training but also potentially lowers the environmental impact. The proposed methods enable researchers with limited resources to participate in large-scale machine learning research, potentially accelerating progress in the field through wider community involvement.
Appendix A Experimental setups
@@
To begin, we introduce the hyperparameter density $\rho$. This hyperparameter represents the fraction of the total space in Linear layers that is updated with a stateful optimizer. For GaLore, this parameter is equal to $\rho=r/h$, where $r$ is the projection rank, and $h$ is the hidden size of the model. For the RandK projection, this parameter can be expressed as $1-s$, where $s$ means sparsity. For BAdam and FRUGAL with the blockwise update, this parameter denotes the ratio of the number of active blocks $a_{\text{block}}$ to the total number of blocks $p$, that is, $\rho=a_{\text{block}}/p$. When using FRUGAL with the column-wise update, as in Section 7, $\rho$ is equal to the ratio of the number of active columns $a_{\text{column}}$ to their total number $h$, i.e., $\rho=a_{\text{column}}/h$.
A.1 Pre-training setup
@@
We adopt a LLaMA-based architecture with RMSNorm (Zhang & Sennrich, 2019) and SwiGLU (Shazeer, 2020) activations on the C4 dataset. Following Zhao et al. (2024a), we trained using a batch size of 512 sequences, sequence length of 256, weight decay of 0, and no gradient clipping. We used T5 tokenizer, since it also was trained on C4 with dictionary size equal to 32k. The update frequency $T$ is set to 200.
Since, unlike GaLore, we consider not only matrix projections, we decided to generalize the concept of rank $r$. Instead, we use density $\rho$, which represents the proportion of Linear layer parameters in the state-full subspace. Thus, for SVD-like projection as in GaLore, the density equals $\rho=r/h$, where $h$ denotes the hidden dimension of the model. We also should point out that similarly to Zhao et al. (2024a), we keep Embeddings, RMSNorms, and Output layer in the state-full subspace throughout the training and don’t reset the optimizer state for them.
@@
We used standard Adam hyperparameters: $\beta_{1}=0.9,\beta_{2}=0.999,\varepsilon=1e-8$. For all methods except GaLore, we selected the learning rate equal to the optimal learning rate for Adam, which we determined through a grid search among values $[1e-4,3e-4,1e-3,3e-3]$. FRUGAL’s learning rate for the state-free optimizer was set equal to that for the state-full optimizer for simplicity and ease of tuning. For a fair comparison with GaLore (Zhao et al., 2024a), we conducted experiments with two learning rate values: 1) the one specified by the authors in the original paper and 2) the optimal learning rate for Adam, as used for other methods. We did this because the learning rate in the original paper could have been optimized for a different number of iterations.
@@
To match the learning rate changes in the first steps of our training with Zhao et al. (2024a), we used a cosine learning rate schedule with restarts, with a warmup of 10% of the steps in a cycle length, and decay of the final learning rate down to 10% of the peak learning rate. To verify that our results are not sensitive to the choice of scheduler, we repeated the experiments for LLaMA-130M with other schedulers. The results for constant with warm-up and cosine (one cycle) with warm-up schedulers can be found in Tables 16 and 16.
For pre-training GPT-2 124M (Radford et al., 2019) we followed the setup described above except for the tokenizer. We utilized the GPT-2 original tokenizer, with 50257 vocabulary size. The results are presented in Table 12.
Table 8: Comparison of validation perplexity and memory estimation for various optimization methods across LLaMA model scales trained on C4 with $\bm{\beta_{2}=0.95}$. We also indicate the additional memory overhead introduced by the optimization algorithm. The values are calculated assuming that each float value occupies 4 bytes (float32). $\rho$ denotes the proportion of the Linear layer parameters in the state-full subspace. Note that Embeddings, RMSNorms, and Output layer are always trained with AdamW.
60M 130M 350M
AdamW 23.51 (0.43G) 18.32 (1.00G) 14.57 (2.74G)
GaLore, $\rho=0.25$ 26.66 (0.30G) 21.03 (0.54G) 16.79 (1.10G)
BAdam, $\rho=0.25$ 25.40 (0.29G) 20.17 (0.52G) 16.54 (1.05G)
FRUGAL, $\rho=0.25$ 24.07 (0.29G) 18.79 (0.52G) 14.96 (1.05G)
FRUGAL, $\rho=0.0$ 24.54 (0.24G) 19.11 (0.37G) 15.20 (0.49G)
Training tokens 20B 20B 24B
Number of iterations 200k 200k 240k
Method 100k iterations
Adam 21.88
GaLore, $\rho=0.25$ 24.19
BAdam, $\rho=0.25$ 25.03
FRUGAL, $\rho=0.25$ 23.17
FRUGAL, $\rho=0.0$ 22.64
Table 10: Perplexity of LLaMA-130M models pre-trained on C4 for 200k steps with different state-free optimizers for FRUGAL.
Method State-free Validation
optimizer perplexity
Adam 18.13
FRUGAL, $\rho=0.25$ signSGD 18.60
FRUGAL, $\rho=0.25$ SGD 19.11
Table 11: Perplexity of LLaMA-130M models pre-trained on C4 with Lion as state-full optimizer for 200k steps.
Method 200k
Adam 18.13
Lion 18.55
GaLore (+ Lion), $\rho=0.25$ 21.65
FRUGAL (+ Lion), $\rho=0.25$ 18.89
Table 12: Validation perplexity of GPT-2 124M model pre-trained on C4 for 200k steps with various optimization methods.
Method Validation perplexity
Adam 21.94
GaLore, $\rho=0.25$ 25.84
BAdam, $\rho=0.25$ 25.43
FRUGAL, $\rho=0.25$ 23.23
FRUGAL, $\rho=0.0$ 25.04
Method Perplexity
Random 18.50
Ascending 18.54
Descending 18.50
Update frequency $T$ Perplexity
10 18.82
20 18.73
50 18.69
100 18.65
200 18.60
500 18.60
1000 18.61
@@
A.2 RoBERTa fine-tuning setup
The batch size and learning rate values used for FRUGAL in the experiments from Table 6 are presented in Table 18. In all experiments, we set the learning rate for the state-free optimizer to $1/10$ of the learning rate of the state-full optimizer. Other hyperparameters, such as scheduler, number of epochs, maximum sequence length, and warmup ratio, were taken from Hu et al. (2021).
@@ A.2 RoBERTa fine-tuning setup
Appendix B Comparison with Concurrent Methods
In this section, we present a comparison with several concurrent works — namely, Fira (Chen et al., 2024a), LDAdam (Robert et al., 2024), and Adamem (Vyas et al., 2024) — which we discovered only during the final stages of our project. Due to this timing, we did not have the opportunity to conduct comprehensive experimental comparisons with them as we did with other baselines in Table 2.
B.1 Algorithmic comparison
We begin by comparing these methods with FRUGAL from an algorithmic perspective.
Adamem.
Similarly to our framework, Vyas et al. (2024) divides the gradient into two parts: the first part is a projection of the gradient onto the top SVD subspace, while the second part is the residual outside this subspace. The first part is used to update momentum, which is then fed into Adafactor’s preconditioner Shazeer & Stern (2018), while the second part is fed directly into a one-sided Adafactor preconditioner Vyas et al. (2024). The outputs of these preconditioners are then used for the final parameter update. Although the core idea of splitting the gradient into two orthogonal subspaces is analogous to Algorithm 1, Adamem implements only one possible variant for choosing subspace and for updates of each component. Thus, Adamem represents a special case of FRUGAL: with SVD-based projection, Adafactor with momentum as the state-full optimizer, and one-sided Adafactor as the state-free optimizer.
Fira.
Chen et al. (2024a) also apply the idea of decomposing the full-rank gradient into a low-rank subspace gradient and a residual gradient. Similarly to GaLore, the low-rank part is obtained through SVD and passed through an AdamW step; however, unlike GaLore, the residual part is not discarded, but is also used in the update in an SGD-like format without using additional optimizer state. An important feature that Fira introduces is norm-based scaling: this technique adaptively scales the per-column learning rate for the residual part by $\frac{\|\psi_{t}(R_{t})\|}{\|R_{t}\|}$ where $\psi_{t}$ denotes the AdamW update rule. In addition, for enhanced training stability, Fira proposes replacing standard gradient clipping with a norm-growth limiter, which transforms abrupt gradient spikes into gradual, smooth increases.
Thus, the key difference between Fira and FRUGAL is that the state-free update sophisticatedly utilizes information from the state-full update. However, Fira has several limitations. First, Fira relies on computing SVD for projection onto the low-rank subspace, which introduces additional memory and computational overhead, with the computational burden becoming more pronounced as model size increases (see Sections 4 and C. Moreover, it is not readily obvious how to propose an alternative to SVD to eliminate this overhead, because, for example, block projection from BAdam (Luo et al., 2024) cannot work with norm-based scaling (Chen et al., 2024a) by construction. Second, following GaLore, Fira continues using the old optimizer state after updating the projection, which should be suboptimal (see Appendix D). Furthermore, according to the original paper Chen et al. (2024a), the reason for instability in standard gradient clipping may be precisely the moments of low-rank subspace transitions.
We also would like to note separately that neither Fira (Chen et al., 2024a) nor Adamem (Vyas et al., 2024) provide theoretical convergence proofs. In contrast, we provide a proof that recovers the best known convergence rate under conventional assumptions.
LDAdam.
Notably, unlike Fira and Adamem, Robert et al. (2024) identify the addition of momentum and gradient from different subspaces as mathematically incorrect (similarly to Appendix D). To resolve this inconsistency,  Robert et al. (2024), propose to reproject the previous optimizer state (similarly to Algorithm 1). Interestingly, LDAdam differs from FRUGAL, Adamem, and Fira in that each individual step remains low-rank, but unlike GaLore and BAdam, the remaining information is not discarded but preserved in an error feedback buffer, thus allowing LDAdam to have convergence guarantees and demonstrate high performance.
However, LDAdam also has a drawback. The method requires updating the projection at every step; and although Robert et al. (2024) proposed replacing the expensive SVD decomposition with significantly cheaper block power iteration, the necessity of performing this operation at every step causes a slowdown of $\geq 15\%$.
B.2 Experimental comparison
We also present a preliminary experimental comparison of FRUGAL with Adamem (Vyas et al., 2024), Fira (Chen et al., 2024a), and LDAdam (Robert et al., 2024).
Adamem.
We compared FRUGAL with our reimplementation of Adamem (Vyas et al., 2024) on pre-training LLaMA model of 3 sizes — 60M, 130M, and 350M — in the same setup as in Sections 6.1 and 2. Unique Adamem hyperparameters were taken from Vyas et al. (2024). The results presented in Table 20 show that while Adamem performs significantly better than GaLore due to utilizing information from the gradient residual, it still falls slightly short of FRUGAL. However, at this point, we are not ready to claim whether this is a consequence of a less favorable choice of state-full and state-free update rules or potentially suboptimal hyperparameter settings for our experimental setup.
Table 20: Comparison of validation perplexity for AdamW, FRUGAL and AdaMeM across LLaMA model scales trained on C4.
60M 130M 350M
AdamW 22.73 18.13 14.43
AdaMeM, $\rho=0.25$ 23.81 18.99 15.10
FRUGAL, $\rho=0.25$ 23.59 18.60 14.79
FRUGAL, $\rho=0.0$ 24.06 18.90 15.03
Training tokens 20B 20B 24B
Number of iterations 200k 200k 240k
Fira and LDAdam
For experiments with Fira (Chen et al., 2024a) and LDAdam (Robert et al., 2024), we deviated from the setup in Appendices A and 2. This modification was necessary since our main setup following GaLore (Zhao et al., 2024a) does not use gradient clipping, while its counterpart—the norm-growth limiter—is critically important for Fira’s performance. Additionally, we observed that Fira significantly benefits from using weight decay, so for the experiments in Table 21 we enabled both gradient clipping (norm-growth limiter for Fira) and weight decay. The results show that all three methods achieve performance very close to AdamW. Surprisingly, LDAdam even outperforms the full-rank AdamW baseline for the 350M model (which, however, may be a result of insufficient hyperparameter grid search). Fira performs slightly better than FRUGAL.
However, we would like to note that these improvements are valid in terms of the number of iterations, and not in terms of wall clock time, where both Fira and LDAdam introduce noticeable time overhead. In the same table, we report the approximate wall-clock slowdown of the methods compared to plain AdamW. We provide approximate values taken from the original papers because the actual slowdown can vary significantly depending on the setup (number of workers in training, GPU type, etc.). One can observe that the slowdown due to the need to perform additional heavyweight operations during the optimizer step can negate the advantages gained from a more carefully designed update rule. Thus, practitioners should choose the algorithm for their specific use case.
Table 21: Comparison of validation perplexity for AdamW, FRUGAL  Fira and LDAdam across LLaMA model scales trained on C4. We also report approximate slowdown comparing to AdamW (taken from original papers).
Approximate Slowdown 130M 350M
AdamW - 17.52 13.81
Fira, $\rho=0.25$ $10\%$ 17.58 13.74
LDAdam, $\rho=0.25$ $15\%$ 17.54 13.54
FRUGAL, $\rho=0.25$ 0% 17.58 13.99
Appendix C Memory estimation